Assume |VBE|= 0.7 V, β=99, VT=25mV. Complete the DC analysis for the circuit (Compute all branch currents and terminal voltages).Problem1: Assume |VE|= 0.7 V, B=99, VT=25mV. Complete the DC analysis for the circuit (Compute all branch currents and termin
Answer #1

Colution Guiven, [VIBE l = 0.7V, B=99, Vr= 25mv. - For F DC- Analysis, All ac-sources are removed T lov All Coupling CapacitoVBI Rth = (R111R2) Rth= (30k 117.5 k Thevenin Impedance acron R2 Tam Rih & With = R2 (1o) = 705 (10) RitR2 30+7.5 [1th = 2v (